OpenStudy (anonymous):

Change the given equation into slope-intercept form. 2x – 5y = -10

OpenStudy (angela210793):

-5y=-10-2x y=2+2x/5

OpenStudy (saifoo.khan):

\[y = \frac25x+2 \]